OpenStudy (delirium):

How will changing electric charge change the equipotential at a distance of 2 meters on an electric field? Medal and fan for help, as always. ;)

OpenStudy (delirium):

@Mehek14 @Conqueror @King.Void. @Awolflover1

OpenStudy (delirium):

Please help.

OpenStudy (delirium):

I don't know what do with that extra bit of information about "2 meters" lol. Increasing the charge increases the equipotential. Decreasing the charge decreases the potential. Distance reduces the intensity of the charge. But I feel like they want a more specific answer...

OpenStudy (samigupta8):

It would simply be kq/r.

OpenStudy (irishboy123):

\(\dfrac{\partial }{\partial q}\left(\dfrac{kq}{r^2}\right)\)